OpenStudy (anonymous):

i got to this part but now i am stuck, can someone please help me? i had to rewrite it in general form. y = 3x^2 – 24x + 10 y – 10 = 3x^2 – 24x y – 10 = 3(x^2 – 8x) y – 10 + 48 = 3(x^2 – 8x + 16)

OpenStudy (anonymous):

@terenzreignz @GoldPhenoix

jimthompson5910 (jim_thompson5910):

y – 10 + 48 = 3(x^2 – 8x + 16) y + 38 = 3(x^2 – 8x + 16) y + 38 = 3(x - 4)^2 (1/3)*(y + 38) = (x - 4)^2 the polynomial is now of the form 4p(y-k) = (x-h)^2

OpenStudy (anonymous):

thank you so much.

jimthompson5910 (jim_thompson5910):

yw